# Figure 1. Confirmation–Diagnosis–Correction–Redo (CDCR) Pattern. Based on the expected time of user interaction under the CDCR pattern, our model decides whether the agent should proceed autonomously or prompt the user to step in and supervise.:

### v1

@@ 5.2. Time Cost
The system measures cost in terms of user interaction time (seconds), following the Confirmation–Diagnosis–Correction–Redo (CDCR) pattern. We first derive the expected time cost from step $i$ to its next confirmation point $j$. The interval begins at a state $s_{i}=\texttt{correct}$, which corresponds to the last state that the user has inspected and determined to be correct. The system then repeatedly chooses to execute actions until it reaches state $s_{j}$. At this point the system chooses to ask the user to verify the correctness of $s_{j}$.
@@
  • $t_{\text{confirm}}^{i}$: time for the user to confirm whether state $s_{i}$ is correct. Concretely, the user compares the screenshot of state with the agent’s activity list (i.e., the sequence of executed actions) to verify correctness.
@@
The first term corresponds to the time it takes the user to confirm state $s_{j}$. The second term corresponds to the case where the state $s_{j}$ is correct, which occurs with probability $p(s_{j}=\texttt{correct}|s_{i}=\texttt{correct})$. If this happens, then the expected remaining time is recursively defined as the expected time needed to traverse from $j$ to $N$ when the system chooses the next checkpoint $k$ such that this time is minimized (denoted as $\min_{k}T[j,k]$). The third term corresponds to the cases where $s_{j}=\texttt{incorrect}$. If this happens, we sum over all the possible locations of the first error weighted by their probability. For each first incorrect state $s_{m}$, the expected remaining time is composed of four parts: the time needed to diagnose the error, the time needed to correct the mistake, the time needed to redo the incorrect steps, and the expected time needed to go from $m$ to $N$ if the system chooses the next checkpoint $k$ such that this time is minimized.
@@ 5.3. Dynamic Programming Checkpoint Solver
Given the recurrent format of T[i, j], we developed a dynamic programming algorithm to efficiently identify the best verification locations that minimize the total user time, see Figure 4.
The algorithm starts at the end of the sequence, since there is always a verification at the end. It then works backwards populating the values of $T[:,:]$ according to the previously defined formula. [^2] We drop $t_{\text{correct}}^{m}$ (the time needed for the user to instruct the agent on how to fix $a_{m}$) from the dynamic programming update because all errors need to be fixed no matter when verifications happen, so it turns out that dropping this term does not change the checkpoint locations. At each iteration, the algorithm computes the best next checkpoint for each starting location $i$, denoted as next_ckpt[i]next\_ckpt[i].
The populated next_ckptnext\_ckpt vector can be used to determine the best checkpoint locations. For a given starting location $i$ (which assumes $s_{i}=\texttt{correct}$), next_ckpt[i]next\_ckpt[i] specifies the location of the best next checkpoint. For the example in Figure 4, when starting at state 1, the best next checkpoint is at state 2, if starting at state 2, the best next checkpoint is at 3, and finally when starting at state 3, the next best checkpoint is at state 5.
@@
5.4. Modeling Personalization Capability
In our system, we have both agent-side and human-side quantities. On the agent side, evaluating agent performance is an active research area, with benchmarks reporting both task accuracy (Xie et al., 2024a; Valmeekam et al., 2023; Xie et al., 2024b; Zhou et al., 2023; Xu et al., 2024; Ma et al., 2024) and execution time (Zhang et al., 2025b, a; Liu et al., 2023; Waytowich et al., 2024). Accordingly, in our model, the step-level accuracy $p_{a_{i}}$ and the execution time $t_{\text{redo}}^{k}$ can be grounded in these benchmark results. For example, in image editing tasks, benchmarks report the success rate of invoking a given tool as well as the average time required to execute a single operation.
On the human side, our framework allows both population-level and per-user parameterization of timing variables such as $t_{\text{confirm}}^{k}$ and $t_{\text{diagnose}}^{k}$. These quantities can capture hierarchical effects, with individual reaction times treated as adjustments to the population average and refined as more user data become available. In this paper, however, we focus on the population-level model and leave personalization to future work.
@@ 6.1. Simulated Environment
Real AI agents vary widely in execution time and outcome quality. We built a simulated environment that fixed execution time and accuracy for each task, allowing us to isolate confirmation frequency as the primary factor. The user interface used a clickable-node tracking design (shown in Figure 5), which participants in our formative study reported as helpful for navigation and understanding task progress. We selected three representative task domains: shopping cart management, image editing, and the Overcooked game. These cover three broad categories: daily life, work-related, and virtual tasks, while also differing substantially in execution time, tool usage, and workflow complexity. Error cases were drawn from real agent failures observed in prior experiments: in shopping, the agent failed to replace an out-of-stock item; in image editing, the agent applied an incorrect mask that prevented object removal; and in Overcooked, the agent grounded to the wrong ingredient dispenser (e.g., selecting the tomato instead of the onion).
Figure 5. Example of the simulated shopping task interface with intermediate confirmation, the activity list on the left and the current action screenshot on the right, enabling users to compare and identify the first error.
@@ 6.5. Variable Determination
The confirmation frequency in our model is determined by four parameters: $t_{\text{confirm}}$, $t_{\text{diagnose}}$, $t_{\text{redo}}$, and the accuracy rate. In this study, each task domain is associated with its own set of parameters. (1) Confirmation and diagnosis times were estimated from a pilot study with random checkpoints. In the main study, these averages were used as fixed inputs, with recalibration performed every 5–10 participants to prevent drift. We deliberately avoided per-user online adaptation, since stable personalization requires long-horizon usage data beyond the scope of this lab setting. (2) Execution time was estimated through formative agent runs and prior literature (Zhang et al., 2025b, a; Liu et al., 2023; Waytowich et al., 2024). We set $t_{\text{redo}}=20$s for shopping (dominated by network and screenshot analysis), $10$s for image editing (external tool calls), and $10$s for Overcooked (agent reasoning plus game state processing). (3) For accuracy, we set fixed per-step action correctness probabilities for each task domain, derived from formative agent runs: $p_{\text{shopping}}=87.5\%$, $p_{\text{image editing}}=91\%$, and $p_{\text{overcooked}}=93\%$. These values were applied uniformly across all steps within a task.
@@ 7.1. Quantitative Results
Across all task domains, our model reduced the completion time of a full task trial (8–16 steps) by 13.54%, corresponding to a savings of 35.84 seconds per task, compared to the baseline (End Confirmation) condition, $t(143)=5.52,p<0.001$. The baseline condition with end confirmation required an average of 264.70 seconds (95% CI [253.43, 275.98]), while intermediate confirmation using our model required 228.86 seconds (95% CI [219.73, 237.99]). Breaking down by task domain, the reductions were 17.44% (47.60s) for Shopping ($t(47)=5.25,p<0.001$; our model mean = 225.33s, 95% CI [212.78, 237.88]), 7.46% (17.41s) for Image Editing ($t(47)=2.34,p=0.024$; our model mean = 215.85s, 95% CI [203.42, 228.28]), and 15.64% (49.03s) for Overcooked Cooking ($t(40)=2.14,p=0.044$; our model mean = 264.51s, 95% CI [236.71, 292.30]).
@@ 8. Discussion
Our findings suggest that confirmation strategies should not be seen as a binary choice between user-driven supervision and agent-driven autonomy. Our decision-theoretic model strikes a mixed-initiative balance, where responsibility is flexibly shared between human and agent. To move toward this vision, we consider two complementary directions: (1) what else to model—expanding beyond efficiency to capture subjective costs and trust, and (2) how else to design—re-imagining confirmation as an interaction paradigm rather than a discrete interruption.
8.1. Beyond Efficiency: Subjective Costs and Trust
@@
8.1.1. Real-world cost
In our CDCR model, cost in the final redo stage is defined only from the agent’s perspective (i.e., the time required for the agent to recover from an error). In practice, however, human users often bear additional costs that extend far beyond execution time. Errors may lead to financial loss (e.g., rescheduling fees for a wrong flight), emotional or social consequences (e.g., accidentally sending an email to the wrong recipient), or even irreversible risks (e.g., permanently deleting a file). Future research should explore how these human-centered costs affect confirmation strategies. For high-stakes actions, mandatory checkpoints should be enforced regardless of the system’s estimated accuracy. For general actions, the replanning term can be extended into a multi-dimensional cost function that balances a heterogeneous variety of costs, such as time, money, emotional costs, and risks.
Reliability engineering provides inspiration, with established methods for integrating economic trade-offs, workforce allocation, and interdependent subsystems into cost functions (Ten Wolde and Ghobbar, 2013; Iren and Bilgen, 2014; Zhou et al., 2015), which could guide extensions of our model.
8.1.2. Perceived reliability
Users’ willingness to confirm is shaped not only by actual system accuracy but also by how reliable they believe the agent to be, based on their subjective experiences and agent framing (Lee and See, 2004; Parasuraman and Manzey, 2010; Shneiderman, 2020). For instance, positive framing can inflate perceived reliability and encourage overuse even when this harms performance (Cockburn et al., 2020; Quinn and Cockburn, 2020, 2016). Our own studies echoed these findings. First-time users who were unfamiliar with the agent tended to perceive it as less reliable, which naturally led them to confirm more frequently. Conversely, in tasks they considered simple, users expected that fewer confirmations would suffice. Yet this expectation often diverged from reality when the agent’s actual accuracy was lower than anticipated.
Such mismatches highlight the need for balancing actual system accuracy with user perception. Our model is already equipped to support this: by assigning each action its own correctness probability $p_{a_{i}}$ based on past performance data, we can introduce a weighting scheme that blends objective error rates with subjective perceptions, ensuring that confirmation policies remain both efficient and aligned with user expectations.
8.2. Designing Confirmation as Interaction

### v2

@@ 5.2. Time Cost
Figure 4. The Confirmation–Diagnosis–Correction–Redo (CDCR) interaction pattern and how our model schedules user verification. Given a multi-step agent plan, the model selects when the user should step in (purple labels). At each scheduled checkpoint, the user (1) Confirms whether the interval’s output matches expectations; if something looks wrong, the user (2) Diagnoses the interval by inspecting nodes one by one to locate the first erroneous step; the user then (3) Corrects the error; and finally the agent (4) Redoes the remaining steps from the corrected point.
The system measures cost in terms of user interaction time (seconds), following the Confirmation–Diagnosis–Correction–Redo (CDCR) pattern, as shown in figure  4. We first derive the expected time cost from step $i$ to its next confirmation point $j$. The interval begins at a state $s_{i}=\texttt{correct}$, which corresponds to the last state that the user has inspected and determined to be correct. The system then repeatedly chooses to execute actions until it reaches state $s_{j}$. At this point, the system chooses to ask the user to verify the correctness of $s_{j}$.
@@
  • $t_{\text{confirm}}^{i}$: time for the user to confirm whether state $s_{i}$ is correct. Concretely, the user compares the screenshot of the state with the agent’s activity list (i.e., the sequence of executed actions) to verify correctness.
@@
The first term corresponds to the time it takes the user to confirm state $s_{j}$. The second term corresponds to the case where the state $s_{j}$ is correct, which occurs with probability $p(s_{j}=\texttt{correct}|s_{i}=\texttt{correct})$. If this happens, then the expected remaining time is recursively defined as the expected time needed to traverse from $j$ to $N$ when the system chooses the next checkpoint $k$ such that this time is minimized (denoted as $\min_{k}T[j,k]$). The third term corresponds to the cases where $s_{j}=\texttt{incorrect}$. If this happens, we sum over all possible locations of the first error, weighted by their probabilities. For each first incorrect state $s_{m}$, the expected remaining time is composed of four parts: the time needed to diagnose the error, the time needed to correct the mistake, the time needed to redo the incorrect steps, and the expected time needed to go from $m$ to $N$ if the system chooses the next checkpoint $k$ such that this time is minimized.
@@ 5.3. Dynamic Programming Checkpoint Solver
Figure 5. Dynamic programming algorithm and results for an agent action sequence of length $N=5$ with action success probabilities $p=[0.7,0.7,0.9,0.85,0.85]$. Costs are parameterized as $t_{\text{confirm}}[k]=1$, $t_{\text{diagnose}}[k]=1$, and $t_{\text{redo}}[k]=1$, for all $k\in\{0,\dots,N\}$. Each cell $\textit{start},\textit{end}]$ specifies the expected user time to correctly execute the rest of the action sequence from startstart to $N$, if the system has a verification checkpoint at state end. White checkmarks indicate the next_ckptnext\_ckpt locations for each startstart. The right-hand annotations summarize the resulting checkpoint policy.
Given the recurrent format of T[i, j], we developed a dynamic programming algorithm to efficiently identify the best verification locations that minimize the total user time, as shown in Figure  5.
The algorithm starts at the end of the sequence, since there is always a verification at the end. It then works backwards, populating the values of $T[:,:]$ according to the previously defined formula. [^2] We drop $t_{\text{correct}}^{m}$ (the time needed for the user to instruct the agent on how to fix $a_{m}$) from the dynamic programming update because all errors need to be fixed no matter when verifications happen, so it turns out that dropping this term does not change the checkpoint locations. At each iteration, the algorithm computes the best next checkpoint for each starting location $i$, denoted as next_ckpt[i]next\_ckpt[i].
The populated next_ckptnext\_ckpt vector can be used to determine the best checkpoint locations. For a given starting location $i$ (which assumes $s_{i}=\texttt{correct}$), next_ckpt[i]next\_ckpt[i] specifies the location of the best next checkpoint. For the example in Figure 5, when starting at state 1, the best next checkpoint is at state 2, if starting at state 2, the best next checkpoint is at 3, and finally, when starting at state 3, the next best checkpoint is at state 5.
@@
5.4. Parameter Realism
In our system, we have both agent-side (probability and redo time) and human-side parameters (confirmation and diagnosis time). The dynamic programming checkpoint solver itself is agnostic to how these parameters are obtained. Realistic values for these parameters can be calibrated from benchmark logs and user studies, and our algorithm permits these parameters to vary step-by-step, drift over time or update online, as illustrated via simulations (Appendix  C).
5.4.1. Probability
On the agent side, evaluating agent performance is an active research area, with benchmarks reporting task accuracy (Xie et al., 2024b; Valmeekam et al., 2023; Xie et al., 2024a; Zhou et al., 2023; Xu et al., 2024; Ma et al., 2024). From such traces, we can estimate per-step success probabilities $p_{a_{i}}$. Crucially, the model does not require a single constant $p$ across the entire task. We allow each step $i$ to have its own success probability $p_{a_{i}}$, which can depend on the tool, page template, or subtask. Appendix C.1.1 shows how our model schedule checkpoints with heterogeneous step-wise accuracies.
For practical deployments, where general benchmark data may differ from the user’s specific environment, $p_{a_{i}}$ can be dynamically estimated using a Beta–Bernoulli belief model: $p_{a_{i}}\sim\mathrm{Beta}(\alpha_{i},\beta_{i})$. In this setup, accuracy metrics reported in benchmarks serve as informative priors. As environment conditions change (e.g., a website updates its layout or the user provides corrective feedback), the belief parameters can be updated online: incrementing $\alpha_{i}$ for autonomous successes and $\beta_{i}$ for failures. The posterior mean would then serve as the input to our dynamic program. While our core algorithm is agnostic to the source of $p_{a_{i}}$, this Bayesian approach demonstrates how the system could adapt to user-specific reality. Determining the optimal weight of the prior relative to new observations remains a nuanced trade-off for future exploration.
5.4.2. Redo
The execution time $t_{\text{redo}}^{k}$ can be grounded in timing benchmark results (Zhang et al., 2025a, b; Liu et al., 2023; Waytowich et al., 2024). Our model supports online replanning. When an error is confirmed at step $m$, we conceptually “cut” the original plan at that point: all steps up to m1m-1 are now treated as confirmed, and the remaining steps form a residual task of length $K^{\prime}$ with a (possibly) updated action sequence. We then update the step-level parameters for this residual task (e.g., new success probabilities and redo times for the revised actions) and recompute the optimal checkpoint schedule for the remaining $K^{\prime}$ steps using the same algorithm. In other words, the same dynamic program is simply re-invoked on the current residual plan, yielding a receding-horizon policy that naturally accommodates updated plans, changing hazards, and revised redo times without altering the core recursion.
5.4.3. User Timing Parameter
On the user side, our model includes two timing parameters: $t_{\text{confirm}}^{k}$ and $t_{\text{diagnose}}^{k}$. These are per-step quantities that may depend on interval-level or interface-level properties, such as the number of steps since the last checkpoint or the complexity of the UI (e.g., DOM size, density of interactive elements). Appendix C.1.2 presents one concrete example. In practice, these parameters can be treated as functions of such features and estimated from pilot data or interaction logs. While different factors influence the numerical values of the timing parameters—and therefore the resulting optimal schedule—they do not change the structure of our algorithm. A more detailed characterization of how specific interface factors shape these two parameters is an important direction for future work.
These two parameters $t_{\text{confirm}}^{k}$ and $t_{\text{diagnose}}^{k}$ can also be personalized. These quantities can capture hierarchical effects, with individual reaction times treated as adjustments to the population average and refined as more user data becomes available. In this paper, however, we focus on the population-level model and leave personalization to future work.
@@ 6.1. Simulated Environment
Real AI agents vary widely in execution time and outcome quality. We built a simulated environment that fixed execution time and accuracy for each task, allowing us to isolate confirmation frequency as the primary factor. The user interface used a clickable-node tracking design (shown in Figure 6), which participants in our formative study reported as helpful for navigation and understanding task progress. We selected three representative task domains: shopping cart management, image editing, and the Overcooked game. These cover three broad categories: daily life, work-related, and virtual tasks, while also differing substantially in execution time, tool usage, and workflow complexity. Error cases were drawn from real agent failures observed in prior experiments: in shopping, the agent failed to replace an out-of-stock item; in image editing, the agent applied an incorrect mask that prevented object removal; and in Overcooked, the agent grounded to the wrong ingredient dispenser (e.g., selecting the tomato instead of the onion). Appendix  B shows details about the simulation environment.
Figure 6. Diagnosis interface with clickable-node tracking for the Home Office Shopping task. (a) The user’s natural-language request to the agent. (b) Actions that the user has confirmed as correct. (c) The clickable-node tracking panel: each past action appears as a node that the user can click to open the corresponding agent execution state (“Desktop View”) or mark as the first wrong step (“Mark as Wrong”), which flags the error and triggers the agent to redo all subsequent actions. (d) The remote desktop view showing the agent’s execution state for the currently selected node.
@@ 6.2. Within-Subjects Design
6.3. Baseline Setting
We selected confirm-at-the-end as our baseline because it is the dominant confirmation strategy in current agentic systems. As we observed in both system review and formative study in Section 3, current agents typically execute tasks autonomously from start to finish, and users only inspect the final result. Although some systems provide reactive controls, which are entirely user-initiated and offer no guidance on when the agent should proactively request confirmation. Because no deployed system offers such functionality, there is no established design to instantiate. In addition, introducing multiple hand-crafted policies would multiply study conditions without providing a principled comparison. For these reasons, we treat “confirm at the end” as the ecologically valid status quo. To complement the user study, we also provide simulations of several alternative confirmation strategies in Appendix  D, allowing systematic comparison without burdening participants.
@@ 6.6. Variable Determination
The confirmation frequency in our model is determined by four parameters: $t_{\text{confirm}}$, $t_{\text{diagnose}}$, $t_{\text{redo}}$, and the accuracy rate. Although the model is fully capable of supporting heterogeneous, per-step values, in our study, we instantiated these parameters at the per-domain level, reflecting the relative uniformity of actions within each task domain. For instance, shopping consists of repeatedly adding different items to a cart based on a predefined shopping list (e.g., a recipe list or office-supply checklist); image editing involves common editing operations such as object removal or color adjustment; and Overcooked involves primitive moves such as navigating to objects or placing items. Using fixed per-domain parameters avoids the high-variance “lucky vs. unlucky” trajectories that arise when per-step probabilities and timings are randomly varied, which would inject noise unrelated to our research question and obscure differences between confirmation strategies. Thus, instead of mixing heterogeneous values within a single trial and producing unstable or uninterpretable trajectories, we evaluate three domains with distinct parameter profiles to show that our model operates robustly across diverse settings.
Parameter values for the three domains were grounded in pilot studies and prior work. (1) Confirmation and diagnosis times were estimated from a 10-participant pilot study with random checkpoints. In the main study, these averages were used as fixed inputs, with recalibration performed every 5–10 participants to prevent drift. We deliberately avoided per-user online adaptation, since stable personalization requires long-horizon usage data beyond the scope of this lab setting. (2) Execution time was estimated through formative agent runs and prior literature (Zhang et al., 2025a, b; Liu et al., 2023; Waytowich et al., 2024). Since steps are similar within each task, we applied fixed per-domain times: $t_{\text{redo}}=20$s for shopping (dominated by network and screenshot analysis), $10$s for image editing (external tool calls), and $10$s for Overcooked (agent reasoning plus game state processing). (3) For accuracy and steps, we set fixed per-step action correctness probabilities for each task domain, derived from formative agent runs: $p_{\text{shopping}}=87.5\%$ with 8 steps, $p_{\text{image editing}}=91\%$ with 12 steps, and $p_{\text{Overcooked}}=93\%$ with 16 steps. These values were applied uniformly across all steps within a task. To contextualize this parameter regime, a per-step success rate of 0.9 across a 10-step task yields roughly $0.9^{10}\approx 0.35$ end-to-end accuracy, consistent with the $\sim 30\%$ accuracy of recent multi-step benchmarks (Ma et al., 2024; Xie et al., 2024b; Xu et al., 2024).
@@ 7.1. Quantitative Results
Across all task domains, our model reduced the completion time of a full task trial (8–16 steps) by 13.54%, corresponding to a savings of 35.84 seconds per task, compared to the baseline (End Confirmation) condition, $t(143)=5.52,p<0.001$. The baseline condition with end confirmation required an average of 264.70 seconds (95% CI [253.43, 275.98]), while intermediate confirmation using our model required 228.86 seconds (95% CI [219.73, 237.99]). Breaking down by task domain, the reductions were 17.44% (47.60s) for Shopping ($t(47)=5.25,p<0.001$; our model mean = 225.33s, 95% CI [212.78, 237.88]), 7.46% (17.41s) for Image Editing ($t(47)=2.34,p=0.024$; our model mean = 215.85s, 95% CI [203.42, 228.28]), and 15.64% (49.03s) for Overcooked ($t(40)=2.14,p=0.044$; our model mean = 264.51s, 95% CI [236.71, 292.30]).
@@ 8. Discussion
Over the past decades, mixed-initiative systems have been moving from user-led tools (Li et al., 2025a; Arakawa et al., 2024; Pu et al., 2025b) toward increasingly agent-led workflows (OpenAI, 2025; Anthropic, 2024). As Tennenhouse (2000) notes, humans shift from being “in the loop” on every step to acting as supervisors and policy-makers. Yet most current agents are still fundamentally reactive: they either wait for users to notice problems and intervene, or they only ask for input when they cannot proceed on their own. Our work aims at the next step in this trajectory: agents that remain mostly autonomous, but proactively allocate supervision by deciding when users should be brought back into the loop. Our scheduler can be plugged into existing systems as a timing layer that determines when to pause, ask for clarification, or surface plans for review. To move toward this vision, we consider two complementary directions: (1) what else to model—expanding beyond efficiency to capture subjective cost, safety, and trust, and (2) how else to design—reimagining confirmation as an interaction paradigm rather than a discrete interruption.
8.1. Beyond Efficiency: Cost, Safety and Trust
@@
8.1.1. Real-world cost and safety
In practice, agentic failures can have an outsized real-world impact that extends far beyond execution time. Errors may lead to financial loss (e.g., rescheduling fees for a wrong flight), emotional or social consequences (e.g., accidentally sending an email to the wrong recipient), or even irreversible risks (e.g., permanently deleting a file) and harms (e.g., privacy breaches, opening phishing links, manipulative suggestions) (Alberts et al., 2024; Bansal et al., 2024; Amodei et al., 2016). Strategic human oversight is therefore crucial for addressing these safety issues. Since human oversight is expensive, it cannot be applied uniformly to every low-level action; instead, it must be focused where it matters most (Holzinger et al., 2025; Amodei et al., 2016). Our CDCR framework speaks directly to this question of where oversight should be applied. Future work should expand the range of costs that can be supported. For general actions, the replanning term (e.g., $t_{redo})$ can be extended to support multi-dimensional cost functions that balance heterogeneous factors, such as time, money, emotional costs, and risks, replacing the purely time-based redo cost used in our current instantiation. Further, future work should also explore the use of mandatory checkpoints for high-stakes actions that occur regardless of the system’s estimated accuracy.
8.1.2. Trust calibration across accuracy regimes
Users’ willingness to confirm is shaped not only by actual system accuracy but also by how reliable they believe the agent to be, based on subjective experience and agent framing (Lee and See, 2004; Parasuraman and Manzey, 2010; Shneiderman, 2020). Prior work shows that overly positive framings of system performance can inflate perceived reliability and encourage overuse even when this harms task outcomes (Cockburn et al., 2020; Quinn and Cockburn, 2020, 2016). Benchmark evaluations can exacerbate this gap: agents may appear highly capable on curated benchmarks while generalizing poorly to practical real-world tasks (Fodor, 2025). In such settings, optimistic benchmark-based messaging can lead users to over-trust the agent and reduce or even cease verification, reinforcing automation complacency. Our model offers a way to counter this by grounding the per-step correctness probabilities $p_{a_{i}}$ in observational real-world performance data (e.g., deployment logs or field studies) rather than benchmark scores alone. In that case, our model’s recommended checkpoint schedule will be more conservative, supporting healthier user oversight and making benchmark optimism visible. As agent performance improves and true error rates drop, the role of user verification in our model correspondingly shifts: its value is less about catching frequent errors and more about maintaining calibrated trust, ensuring users retain occasional oversight instead of becoming fully complacent.
8.2. Designing Confirmation for Transparency